# Condense, Don’t Just Prune: Enhancing Efficiency and Performance in MoE Layer Pruning

### v2

@@ Abstract
Mixture-of-Experts (MoE) has garnered significant attention for their ability to scale up neural networks while utilizing the same or even fewer active parameters. However, MoE does not relieve the massive memory requirements of networks, which limits their practicality in real-world applications, especially in the era of large language models (LLMs). While recent work explores the possibility of removing entire layers of MoE to reduce memory, the performance degradation is still notable. In this paper, we propose ConDense-MoE (CD-MoE) that, instead of dropping the entire MoE layer, condenses the large, sparse MoE layer into a smaller, denser layer with only a few experts activated for all tokens, while maintaining hardware friendliness. Our approach is specifically designed for fine-grained MoE with shared experts, where Feed-Forward Networks are split into many small experts, with certain experts isolated to serve as shared experts that are always activated, such as DeepSeekMoE and QwenMoE. We demonstrate the effectiveness of our method. Specifically, for the DeepSeekMoE-16B model, our approach maintains 90% of the average accuracy while reducing memory usage by 27.5% and increasing inference speed to 1.26 times. Moreover, we show that by applying lightweight expert fine-tuning—only to the condensed layers—and using 5 hours on a single 80G A100 GPU, we can successfully recover 98% of the original performance. Our code is available at https://github.com/duterscmy/CD-MoE/tree/main.
Condense, Don’t Just Prune:
Enhancing Efficiency and Performance in MoE Layer Pruning
Mingyu Cao1, Gen Li2, Jie Ji2, Jiaqi Zhang3, Xiaolong Ma2, Shiwei Liu4, Lu Yin5thanks: Correspondence to Lu Yin: l.yin@surrey.ac.uk 1Shopee Pte. Ltd, China  2Clemson University, USA  3Meituan Inc, China 4University of Oxford, United Kingdom  5University of Surrey, United Kingdom
@@
1 Introduction
Large Language Models (LLMs) continue to grow in both size and capability (Brown, 2020; Touvron et al., 2023b, a), fueling demand for architectures that can scale with minimal increases in computational cost. Mixture of Experts (MoE) architectures have emerged as a promising solution to this challenge (Shazeer et al., 2017; Jiang et al., 2024a; Team et al., 2024). Unlike standard dense networks, MoEs selectively activate only the most relevant subset of parameters (referred to as “experts”) for a given input, enabling substantial growth in model capacity without a proportional escalation in compute overhead. Recent advances in fine-grained expert segmentation (Dai et al., 2024; Team, 2024) push this concept further by splitting feed-forward layers into many small experts, while designating a small set of shared experts that remain active across all tokens. Despite the clear benefits, MoE architectures still face major deployment barriers due to high memory costs. Storing a large number of experts, even if most remain dormant per token, can be prohibitively expensive in real-world systems. This reality highlights the urgent need for approaches that preserve the performance benefits of MoEs while lowering their memory footprint.
In this paper, we present a new framework called ConDense-MoE (CD-MoE), a novel framework that significantly enhances the efficacy specifically for DeepSeek-style shared-expert-type MoE (Dai et al., 2024; Team, 2024). Unlike prior MoE compression approaches that often remove entire layers or retain the routing mechanism with fewer experts (He et al., 2024), our method eliminates the routing process at selected layers and condenses them into dense layers. Specifically, we devise a greedy strategy to identify which layers—and which experts within them—can be condensed with minimal impact on model accuracy, leading to significant reductions in memory usage and faster inference.
@@ 1 Introduction
  • Empirical validation on DeepSeekMoE-16B, demonstrating a 27.5% memory reduction and a 1.26$\times$ inference speed, while retaining up to 90% of the zero-shot accuracy. Moreover, lightweight fine-tuning can effectively reclaim 98% of the original performance using only a single A100 GPU in a few hours.
Input: Calibration data input $X$, routing expert set $E_{\text{route}}=\{E_{1},\ldots,E_{n}\}$, number of experts to select $K$
Output: Condense expert set $E_{\text{condense}}=\{E_{\text{shared}}\}$
Initialize $E_{\text{condense}}\leftarrow\{E_{\text{shared}}\}$,  $O_{\text{ref}}\leftarrow\text{Layer}_{\text{route}}(X)$trianglerightCompute reference output with all routing experts
for $k\leftarrow 1$ to $K$ do
       for each expert $E_{i}\in E_{\text{route}}$ do
             $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{i}\}$trianglerightTemporarily add expert to condense set
             $O_{\text{tmp}}\leftarrow\text{Layer}_{\text{condense}}(X)$trianglerightCompute output with current condense experts
             $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$trianglerightCompute JS divergence as loss for $E_{i}$
             $E_{\text{condense}}\leftarrow E_{\text{condense}}\setminus\{E_{i}\}$trianglerightRemove expert after computing loss
      $E_{\text{best}}\leftarrow\text{argmin}_{E_{i}\in E_{\text{route}}}\,loss_{i}$trianglerightSelect the expert with minimum loss
       $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{\text{best}}\}$trianglerightAdd best expert to condense set
       $E_{\text{route}}\leftarrow E_{\text{route}}\setminus\{E_{\text{best}}\}$trianglerightRemove selected expert from routing set
Algorithm 1 Condense Experts Selection
@@ 2 Related Work
Mixture-of-Experts (MoE) architectures have become a powerful framework for scaling neural networks without linearly increasing computational burden (Shazeer et al., 2017; Fedus et al., 2022; Jiang et al., 2024a; Dai et al., 2024; Team, 2024). Early methods such as Switch Transformer (Fedus et al., 2022) demonstrated that sparse activation can substantially boost model capacity while maintaining tractable compute. Further work has refined routing policies (Jiang et al., 2024b) and introduced a fine-grained segmentation of feed-forward layers into many small experts, often augmented with “shared” experts that remain active for every token (Dai et al., 2024; Team, 2024). These advances collectively enhance training stability and generalization, yet the growing pool of experts poses serious deployment challenges, chiefly due to large memory demands.
A range of pruning and compression strategies have emerged to address these overheads. Some works prune experts by identifying those that are minimally activated or less essential for downstream tasks (Chi et al., 2022; Sun et al., 2024; Muzio et al., 2024), with solutions ranging from exhaustive searches over all expert subsets (Lu et al., 2024) to filtering experts by their activation frequency. While these methods typically preserve routing, more aggressive approaches prune entire MoE layers, achieving considerable memory reductions but at the risk of pronounced accuracy losses (He et al., 2024). In contrast, our technique selectively eliminates the routing mechanism in particular layers and retains only a small subset of experts (including the shared experts), thus reducing memory demands while sustaining most of the model’s representational power.
@@ 3 ConDense-MoE (CD-MoE)
Input: Data input $X$, routing layer set $L_{\text{route}}=\{L_{1},\ldots,L_{n}\}$, number of layers to select $K$
Output: Condense layer set $L_{\text{condense}}=\{\}$
Initialize $L_{\text{condense}}\leftarrow\{\}$,  $O_{\text{ref}}\leftarrow\text{Model}(X)$trianglerightCompute reference output with all routing layers
for $k\leftarrow 1$ to $K$ do
       for each layer $L_{i}\in L_{\text{route}}$ do
             $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{i}\}$trianglerightTemporarily add layer to condense set
             $O_{\text{tmp}}\leftarrow\text{Model}(X)$trianglerightCompute output with current condense layers
             $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$trianglerightCompute JS divergence as loss for $L_{i}$
             $L_{\text{condense}}\leftarrow L_{\text{condense}}\setminus\{L_{i}\}$trianglerightRemove layer after computing loss
      $L_{\text{best}}\leftarrow\text{argmin}_{L_{i}\in L_{\text{route}}}\,loss_{i}$trianglerightSelect the layer with minimum loss
       $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{\text{best}}\}$trianglerightAdd best layer to condense set
       $L_{\text{route}}\leftarrow L_{\text{route}}\setminus\{L_{\text{best}}\}$trianglerightRemove selected layer from routing set
Algorithm 2 Condense Layers Selection
@@ 3.1 Preliminaries of MoE
$g_{i,t}=\begin{cases}s_{i,t},&s_{i,t}\in\text{TopK}\left(\{s_{j,t}\mid 1\leq j\leq N\},K\right),\\ 0,&\text{otherwise},\end{cases}$
@@ 3.2 Expert Selection and Condensing
Our CD-MoE framework removes the standard routing mechanism during inference. For each expert $E_{i}$, we use C4 calibration data to inference and compute the mean of $g_{i,t}$ when that expert is activated, which is then used as the fixed gate value $g_{i}$ for that expert during the inference phase:
$g_{i}=\frac{1}{|\{t\mid g_{i,t}\neq 0\}|}\sum_{t\mid g_{i,t}\neq 0}g_{i,t}$ (2)
@@
In which, ${JS}(.)$ is the Jensen-Shannon divergence used to measure the difference between two outputs. Zhang et al. (2024) shown that in model pruning, this metric is better than angular distance and Euclidean distance. We also demonstrated its superiority over other metrics such as Kullback–Leibler (KL) divergence and perplexity in Table 2. Therefore, we choose ${JS}(.)$ divergence as our selection metric. The mathematical expression of ${JS}(.)$ is as follows:
$JS(u,v)=\frac{1}{2}\left(KL\left(u\,,\,\frac{1}{2}(u+v)\right)\right.\\ \left.+KL\left(v\,,\,\frac{1}{2}(u+v)\right)\right)$ (4)
Here, ${KL}(.)$ represents the Kullback–Leibler divergence, defined for discrete random variables as:
$KL(u,v)=\sum_{j}u_{j}\log\left(\frac{u_{j}}{v_{j}}\right)$ (5)
where $u$ and $v$ are vectors that represent discrete probability distributions, which, in our method, can be the dense layer output and the condensed layer output.
@@ 3.3 Layer Selection and Condensing
Figure 2: Fluctuations in the $JS$ divergence between the the outputs of the condensed model and the original dense model across different layers.
To systematically quantify the impact of condensing individual layers, we conducted preliminary experiments assessing the output changes using the Jensen-Shannon ($JS$) divergence between the pruned and unpruned model outputs as our evaluation metric. As illustrated in Figure 2, condensing layer 0 results in a significant shift in the output distribution, indicating its critical role in the model’s performance. While the middle layers exhibit some fluctuation, condensing the last three layers also causes noticeable deviations, suggesting that these layers are essential for maintaining output fidelity. These observations indicate that condensing different layers introduces varying degrees of change in the $JS$ metric, reflecting their varying importance to the overall model performance.
Table 1: Comparison with the dense models, unpruned MOE model and the baseline methods, the underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#$\#L$’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model. CD-MoE-S refers to retaining only the Shared expert after condensing, while CD-MoE-SR means selecting additional 6 Routed experts and keeping as dense.
@@
CD-MoE-S + SFT 8 2.4B 73.1% 1.34x 41.3 72.1 65.9 39.7 40.0 76.2 67.9 69.7 59.1
CD-MoE-SR + SFT 9 2.8B 72.5% 1.26x 42.6 79.2 66.9 38.6 38.6 75.7 66.8 75.2 60.5
Motivated by these findings, we adopt a greedy search strategy analogous to that used in expert selection to determine the optimal layers for condensing. Specifically, we compute the JS divergence before and after condensing based on the output of the final layer, iteratively adding layers to the $Layer_{condense}$ set that can minimize the distance between the model’s final layer output and the output before condensing, as detailed in Algorithm 2. For layers that are not condensed, the standard token routing mechanisms are employed during forward propagation to maintain their full expressive power. For the layers being condensed, the routing mechanism is removed, and the MOE is compressed into a dense structure. This selective condensing approach allows us to effectively reduce computational complexity and memory usage while minimizing degradation in model performance.
4 Experiments
4.1 Experiment Setup
We adopt the opensource model DeepSeekMoE-16B and Qwen1.5-MoE-A2.7B (Team, 2024). Besides, in alignment with the methodologies outlined by Sun et al. (2024) and Lu et al. (2024), we randomly sample 100 examples from the C4 dataset to serve as calibration data during the condensing phase of our model.
Evaluation: To provide a comprehensive assessment of our model’s performance, we follow the protocol established by He et al. (2024). We report zero-shot accuracies on eight diverse tasks selected from the EleutherAI Language Model Evaluation Harness (Gao et al., 2024), which includes the 7 commonsense reasoning benchmark datasets: ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019a), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), RTE (Dagan et al., 2005), WinoGrande (Sakaguchi et al., 2019), and a more challenging dataset MMLU (Hendrycks et al., 2021).
Baseline: For comparative analysis, we primarily benchmark our approach against the Block Trimming and Layer Trimming methods proposed by He et al. (2024). These methods serve as strong baselines for pruning MoE models and are directly relevant to our study. Both pruning methods prune all experts in the layer. They adopt Block Influence for pruning as introduced in Men et al. (2024).
We also compared the results with dense models that have similar activations, such as GPT-Neo-2.7B (Gao et al., 2020), OPT-2.7B (Zhang et al., 2022), BLOOM-3B (Press et al., 2022), and OpenLLaMA-3B (Touvron et al., 2023a).
@@ 4.1 Experiment Setup
Figure 3: CD-MoE against baselines on zero-shot tasks w/o fine-tuning. Left: Average accuracy with varying Memory Ratio against the original model. Right: Average accuracy with varying SpeedUp against the original model. The Gray dotted line is the original model result. CD-MoE-S represents the shared experts and no routing experts, and CD-MoE-SR represents shared with routing experts. Baseline indicated performance of the dense model.

### v3

@@ Abstract
Mixture-of-Experts (MoE) has garnered significant attention for its ability to scale up neural networks while utilizing the same or even fewer active parameters. However, MoE does not relieve the massive memory requirements of networks, which limits their practicality in real-world applications, especially in the era of large language models (LLMs). While recent work explores the possibility of removing entire layers of MoE to reduce memory, the performance degradation is still notable. In this paper, we propose ConDense-MoE (CD-MoE) that, instead of dropping the entire MoE layer, condenses the large, sparse MoE layer into a smaller, denser layer with only a few experts activated for all tokens, while maintaining hardware friendliness. Our approach is specifically designed for fine-grained MoE with shared experts, where Feed-Forward Networks are split into many small experts, with certain experts isolated to serve as shared experts that are always activated, such as DeepSeekMoE and QwenMoE. We demonstrate the effectiveness of our method. Specifically, for the DeepSeekMoE-16B model, our approach maintains 90% of the average accuracy while reducing memory usage by 27.5% and increasing inference speed to 1.26 times. Moreover, we show that by applying lightweight expert fine-tuning—only to the condensed layers—and using 5 hours on a single 80G A100 GPU, we can successfully recover 98% of the original performance. Our code is available at: https://github.com/duterscmy/CD-MoE.
1 Introduction
Large Language Models (LLMs) continue to grow in both size and capability (Brown, 2020; Touvron et al., 2023b; a), fueling demand for architectures that can scale with minimal increases in computational cost. Mixture of Experts (MoE) architectures have emerged as a promising solution to this challenge (Shazeer et al., 2017; Jiang et al., 2024; Team et al., 2024). Unlike standard dense networks, MoEs selectively activate only the most relevant subset of parameters (referred to as “experts”) for a given input, enabling substantial growth in model capacity without a proportional escalation in compute overhead. Recent advances in fine-grained expert segmentation (Dai et al., 2024; Team, 2024) push this concept further by splitting feed-forward layers into many small experts, while designating a small set of shared experts that remain active across all tokens. Despite the clear benefits, MoE architectures still face major deployment barriers due to high memory costs. Storing a large number of experts, even if most remain dormant per token, can be prohibitively expensive in real-world systems. This reality highlights the urgent need for approaches that preserve the performance benefits of MoEs while lowering their memory footprint.
In this paper, we present a new framework called ConDense-MoE (CD-MoE), a novel framework that significantly enhances the efficacy specifically for DeepSeek-style shared-expert-type MoE (Dai et al., 2024; Team, 2024). Unlike prior MoE compression approaches that often remove entire layers or retain the routing mechanism with fewer experts (He et al., 2024b), our method eliminates the routing process at selected layers and condenses them into dense layers. Specifically, we devise a greedy strategy to identify which layers—and which experts within them—can be condensed with minimal impact on model accuracy, leading to significant reductions in memory usage and faster inference.
@@ 1 Introduction
[…]
@@ 2 Related Work
Mixture-of-Experts (MoE) architectures have become a powerful framework for scaling neural networks without linearly increasing computational burden (Shazeer et al., 2017; Jiang et al., 2024; Dai et al., 2024; Team, 2024). Early methods such as Switch Transformer (Fedus et al., 2022) demonstrated that sparse activation can substantially boost model capacity while maintaining tractable compute. Further work has refined routing policies (Jiang et al., 2024) and introduced a fine-grained segmentation of feed-forward layers into many small experts, often augmented with “shared” experts that remain active for every token (Dai et al., 2024; Team, 2024). These advances collectively enhance training stability and generalization, yet the growing pool of experts poses serious deployment challenges, chiefly due to large memory demands.
A range of pruning and compression strategies have emerged to address these overheads. Some works prune experts by identifying those that are minimally activated or less essential for downstream tasks (Chi et al., 2022; Sun et al., 2024; Muzio et al., 2024), with solutions ranging from exhaustive searches over all expert subsets (Lu et al., 2024) to filtering experts by their activation frequency. While these methods typically preserve routing, more aggressive approaches prune entire MoE layers, achieving considerable memory reductions but at the risk of pronounced accuracy losses (He et al., 2024b). In contrast, our technique selectively eliminates the routing mechanism in particular layers and retains only a small subset of experts (including the shared experts), thus reducing memory demands while sustaining most of the model’s representational power.
@@ 3.1 Preliminaries of MoE
$\small g_{i,t}=\begin{cases}s_{i,t},&s_{i,t}\in\text{TopK}\left(\{s_{j,t}\mid 1\leq j\leq N\},k\right),\\ 0,&\text{otherwise},\end{cases}$
@@ 3.2 Expert Selection and Condensing
Input: Calibration data input $X$, routing expert set $E_{\text{route}}=\{E_{1},\ldots,E_{n}\}$, number of selected experts $K$
Output: Condense expert set $E_{\text{condense}}=\{E_{\text{shared}}\}$
Initialize $E_{\text{condense}}\leftarrow\{E_{\text{shared}}\}$,    $O_{\text{ref}}\leftarrow\text{Layer}_{\text{route}}(X)$ $\triangleright$ //Compute reference output with all routing experts
for $k\leftarrow 1$ to $K$ do
  for each expert $E_{i}\in E_{\text{route}}$ do
     $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{i}\}$ $\triangleright$ //Temporarily add expert to condense set
     $O_{\text{tmp}}\leftarrow\text{Layer}_{\text{condense}}(X)$ $\triangleright$ //Compute output with current condense experts
     $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$ $\triangleright$ //Compute JS divergence as loss for $E_{i}$
     $E_{\text{condense}}\leftarrow E_{\text{condense}}\setminus\{E_{i}\}$ $\triangleright$ //Remove expert after computing loss
 $E_{\text{best}}\leftarrow\text{argmin}_{E_{i}\in E_{\text{route}}}\,loss_{i}$ $\triangleright$ //Select the expert with minimum loss
  $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{\text{best}}\}$ $\triangleright$ //Add best expert to condense set
  $E_{\text{route}}\leftarrow E_{\text{route}}\setminus\{E_{\text{best}}\}$ $\triangleright$ //Remove selected expert from routing set
Algorithm 1 Condense Experts Selection
Our CD-MoE framework removes the standard routing mechanism during inference. For each expert $E_{i}$, we use open-source C4 calibration data to compute the mean of $g_{i,t}$ when that expert is activated, which is then used as the fixed gate value $g_{i}$ for that expert during the inference phase:
$g_{i}=(1/|{t\mid g_{i,t}\neq 0}|)\sum_{t\mid g_{i,t}\neq 0}g_{i,t}$ (2)
@@
In which, $K$ is consistent with the number of routing experts(non-shared experts) activated in the original model, and JS(.){JS}(.) is the Jensen-Shannon divergence used to measure the difference between two outputs. Zhang et al. (2024) shown that in model pruning, this metric is better than angular distance and Euclidean distance. We also demonstrated its superiority over other metrics such as Kullback–Leibler ($KL$) divergence and perplexity in Table 8.
@@ 3.3 Layer Selection and Condensing
Input: Data input $X$, routing layer set $L_{\text{route}}=\{L_{1},\ldots,L_{n}\}$, number of condensed layers NN
Output: Condense layer set $L_{\text{condense}}=\{\}$
Initialize $L_{\text{condense}}\leftarrow\{\}$, $O_{\text{ref}}\leftarrow\text{Model}(X)$ $\triangleright$ //Compute reference output with all routing layers
for $n\leftarrow 1$ to NN do
  for each layer $L_{i}\in L_{\text{route}}$ do
     $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{i}\}$ $\triangleright$ //Temporarily add layer to condense set
     $O_{\text{tmp}}\leftarrow\text{Model}(X)$ $\triangleright$ //Compute output with current condense layers
     $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$ $\triangleright$ //Compute JS divergence as loss for $L_{i}$
     $L_{\text{condense}}\leftarrow L_{\text{condense}}\setminus\{L_{i}\}$ $\triangleright$ //Remove layer after computing loss
 $L_{\text{best}}\leftarrow\text{argmin}_{L_{i}\in L_{\text{route}}}\,loss_{i}$ $\triangleright$ //Select the layer with minimum loss
  $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{\text{best}}\}$ $\triangleright$ //Add best layer to condense set
  $L_{\text{route}}\leftarrow L_{\text{route}}\setminus\{L_{\text{best}}\}$ $\triangleright$ //Remove selected layer from routing set
Algorithm 2 Condense Layers Selection
@@
Figure 2: Left: Fluctuations in the $JS$ divergence between the the outputs of the condensed model and the original dense model across different layers. Middle: fluctuations in the KLKL divergence. Right: fluctuations in the perplexity.
To systematically quantify the impact of condensing individual layers, we conducted preliminary experiments assessing the output changes using the $JS$ divergence, $KL$ divergence and perplexity between the pruned and unpruned model outputs as our evaluation metric. As illustrated in Figure 2, condensing layer 0 results in a significant shift in the output distribution, indicating its critical role in the model’s performance. While the middle layers exhibit some fluctuation, condensing the last three layers also causes noticeable deviations, suggesting that these layers are essential for maintaining output fidelity. These observations indicate that condensing different layers introduces varying degrees of change in the three metrics, reflecting their varying importance to the overall model performance.
Motivated by these findings, we adopt a greedy search strategy analogous to that used in expert selection to determine the optimal layers for condensing. Specifically, we compute the $JS$ divergence before and after condensing based on the output of the final layer, iteratively adding layers to the $Layer_{condense}$ set that can minimize the distance between the model’s final layer output and the output before condensing, as detailed in Algorithm 2. For layers that are not condensed, the standard token routing mechanisms are employed during forward propagation to maintain their full expressive power. For the layers being condensed, the routing mechanism is removed, and the MOE is compressed into a dense structure. The number of condensing layers N in can be chosen based on goals (e.g., memory budget vs. desired accuracy). This selective condensing approach allows us to effectively reduce computational complexity and memory usage while minimizing degradation in model performance.
3.4 Time Complexity Analysis
The computational efficiency of our greedy search algorithms stems from their polynomial time complexity relative to key model parameters:
  • Greedy Expert Selection: For a layer with $E$ routing experts and target selection of $K$ routing experts, each iteration evaluates all remaining candidates . The total complexity is $O(K\cdot E)$, which improves upon exhaustive search’s $O(2^{E})$ complexity (Lu et al., 2024).
  • Greedy Layer Selection: With $L$ candidate layers and $N$ target condensed layers, each step evaluates up to $L$ layers. The complexity becomes $O(N\cdot L)$.
Taking DeepSeekMoE-16B ($E=66$, $L=27$) as an example, selecting $K=6$ routing experts per condensed layer requires only 369 inferences, demonstrating practical scalability. The linear dependence on $E$ and $L$ makes our approach feasible for billion-parameter MoEs.
4 Experiments
4.1 Experiment Setup
Table 1: Results on DeepSeekMoE-16B. The underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#L\#L’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model. The lightweight SFT results are averaged over three runs with different random seeds, and the standard deviation is indicated below the accuracy.
@@ 4.1 Experiment Setup
LayerDrop 8 2.4B 72.2% 1.34x 37.8 71.0 58.9 39.0 34.2 64.5 60.8 70.0 54.5
$\pm$0.7 $\pm$1.0 $\pm$0.4 $\pm$0.7 $\pm$0.9 $\pm$1.1 $\pm$0.4 $\pm$0.3 $\pm$0.3
CD-MoE-S 8 2.4B 73.1% 1.34x 41.5 72.0 66.0 39.5 40.3 75.9 67.7 70.0 59.1
$\pm$0.7 $\pm$0.5 $\pm$0.5 $\pm$0.8 $\pm$0.5 $\pm$0.7 $\pm$1.1 $\pm$0.4 $\pm$0.2
CD-MoE-SR 9 2.8B 72.5% 1.26x 42.4 79.4 66.7 38.5 38.6 76.1 67.0 73.9 60.4
$\pm$0.6 $\pm$0.8 $\pm$0.3 $\pm$0.7 $\pm$0.8 $\pm$0.7 $\pm$0.3 $\pm$1.5 $\pm$0.2
@@
Figure 3: CD-MoE against baselines on zero-shot tasks w/o fine-tuning on DeepSeekMoE-16B (top) and Qwen1.5-MoE-A2.7B (bottom). Left: Average accuracy with varying Memory Ratio against the original model. Right: Average accuracy with varying SpeedUp against the original model. The gray dotted line is the original model result.
We adopt the open-source models DeepSeekMoE-16B and Qwen1.5-MoE-A2.7B (Team, 2024) (with 16B and 14.3B parameters, respectively) as representative high-capacity examples, and a larger model Qwen2-MoE-57B-A14B. We randomly sample 100 examples from the open-source C4 dataset to serve as calibration data during the condensing phase of our model, due to its widespread adoption in the LLM community (Sun et al., 2024; Frantar and Alistarh, 2023; Yin et al., 2024).
Evaluation: To provide a comprehensive assessment of our model’s performance, we follow the protocol established by He et al. (2024b). We report zero-shot accuracies on eight diverse tasks selected from the EleutherAI Language Model Evaluation Harness (Gao et al., 2024), which includes the 7 open-source commonsense reasoning benchmark datasets: ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019a), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), RTE (Dagan et al., 2005), WinoGrande (Sakaguchi et al., 2019), and a more challenging dataset MMLU (Hendrycks et al., 2021). In addition, we evaluate the reasoning ability of the models on the mathematical task GSM8K (Cobbe et al., 2021) and the code generation task HumanEval (Chen et al., 2021).
For quality assessment, we use accuracy as the evaluation metric for commonsense reasoning, MMLU, and GSM8K. For HumanEval, we adopt Pass@1 as the metric. To provide a more detailed comparison across different pruning methods, we employ the following two metrics:
  • SpeedUp: We perform inference 5,000 times on a fixed input sequence and measure the average latency. SpeedUp is computed as $\text{SpeedUp}=\frac{\text{Latency}(\text{Original Model})}{\text{Latency}(\text{Pruned Model})}$.
  • Memory Ratio: We measure the peak GPU memory consumption during inference for the full model and each pruned variant using a fixed input. Memory Ratio is computed as $\text{Memory Ratio}=\frac{\text{Peak Memory}(\text{Pruned Model})}{\text{Peak Memory}(\text{Original Model})}$.
Additionally, the Recovery Ratio reported in our abstract and experiment analysis is computed as: $\frac{\text{Average Accuracy}(\text{Pruned Model})}{\text{Average Accuracy}(\text{Original Model})}$.
Experimental Methods: We primarily compare our method against the Block Drop and Layer Drop (He et al., 2024b) and M-SMoE (Li et al., 2024b). These methods serve as strong baselines for pruning MoE models and are directly relevant to our study. Block Drop and Layer Drop prune all experts and the routing process in the layer, which is more aggressive than our method, while M-SMoE more conservatively retains the routing process and only merges the experts into a smaller number. We also compared the results with dense models that have similar activations, such as GPT-Neo-2.7B (Gao et al., 2020), OPT-2.7B (Zhang et al., 2022), BLOOM-3B (Press et al., 2022), and OpenLLaMA-3B (Touvron et al., 2023a).
For our proposed method, we introduce two variants::
  • CD-MoE-S: This variant retains only the Shared experts after condensation. The number of shared experts varies across different models, as detailed in Appendix A.1.
  • CD-MoE-SR: This variant selects not only the Shared expert(s) but also an additional $K$ Routed experts, preserving them as dense. Here, $K$ is consistent with the number of routed experts activated in the original model, i.e., $K=\text{Number of Activated Experts}-\text{Number of Shared Experts}$ (see Appendix A.1).